OpenStudy (anonymous):

the measure of an angle is 30 more than 4 times the measure of its complement. what are the measures of the angles?

OpenStudy (mathstudent55):

Complementary angles are two angles whose measures add up tp 90. Let the complement be x; 30 more than 4 times the complement then is 30 + 4x; When you add them, you get 90: x + 30 + 4x = 90 Solve for x. That'll be the complement. To get the first angle substitute the value of x in 30 + 4x, and evaluate the expression.

OpenStudy (anonymous):

78 and 12 degrees, Soln: Let one angle be A Complementary angle = 90-A Therefore, A = 30 +4(90-A) A = 30 + 360-4A 5A = 390 A=78 and the other angle => 90-A = 90-78=12

OpenStudy (anonymous):

how about is : the sum of the measures of the interior angles of a polygon is 540 find out the number of sides the polygon has?

OpenStudy (anonymous):

The complement of an angle \(\theta\) is \(90^\circ-\theta\); we can write our description of the measure of the angle mathematically to yield \(\theta=30^\circ+4(90^\circ-\theta)\). We can solve for \(\theta\) as follows:$$\theta=30^\circ+360^\circ-4\theta\\5\theta=390^\circ\\\theta=78^\circ$$Since we know our angle \(\theta\) is \(78^\circ\), its ccomplementary angle is \(90^\circ-\theta=90^\circ-78^\circ=12^\circ\).

OpenStudy (anonymous):

http://www.mathsisfun.com/geometry/interior-angles-polygons.html So we know there's a formula for the sum of our interior angles, \(\Theta=(n-2)\times180^\circ\). We're told the total sum \(\Theta\) is \(540^\circ\). Can you solve for \(n\)?

OpenStudy (mathstudent55):

how about is : the sum of the measures of the interior angles of a polygon is 540 find out the number of sides the polygon has? For this problem, you need to know the following formula: The sum of the measures of the interior angles of an n-sided polygon is (n - 2)180 Since you know the sum is 540, just let 540 equal the formula and solve for n: (n - 2)180 = 540 To solve for n, start by dividing both sides by 540. Then, add 2 to both sides.

OpenStudy (anonymous):

one last problem!! how do you find measure of angles R